An astronaut's pack weighs 14.9 N when she is on earth but only 1.55 N when she is at the surface of moon.

What is the acceleration due to gravity on this moon?

What is the mass of the pack on this moon?
Answer #1

14.9 = m*9.8

m = 1.52 - mass of the pack

1.55 = mg

g = 1.55/1.52 = 1.02 m/s2 - acceleration due to gravity